a collection of dimes and nickels contains 37 coins and has a total value of $2.85.how many of each coin are there in the collection?
Answer by checkley77(12844)

N+D=37 OR N=37-D
.05N+.10D=2.85
NOW REPLACE N IN THE SECOND EQUATION WITH (37-D) & SOLVE FOR THR NUMBER OF DIMES.
.05(37-D)+.10D=2.85
1.85-.05D+.10D=2.85
.05D=2.85-1.85
.05D=1.00
D=1.00/.05
D=20 NUMBER OF DIMES.
N=37-20=17 IS THE NUMBER OF NICKELS.
PROOF:
.05*17+.10*20=2.85
.85+2.00=2.85
5.85=5.85